Factor Completely (difference of squares)
x^8-1

Answer by kapilsinghi(68)

x^8 - 1
=(x^2)^3 - 1^3
applying the formula for a^3 - b^3
=(x^2 - 1)((x^2)^2 + 1 + x^2)
=(x^2 - 1^2)((x^2)^2 + 1 + x^2)
=(x-1)(x+1)(x^2)^2 + 1 + x^2)
kapilsinghi123@gmail.com